OpenStudy (anonymous):

What is the equation of a line parallel to the line y = 4x + 7 and passes through the point (3, 0)?

OpenStudy (anonymous):

A. y = -4x + 12 B. y = 4x − 12 C. y = 4x − 3 D. y = 4x + 3

OpenStudy (anonymous):

B!